### v4

@@ 4.1 Overall structure
Figure 9: Backbone Network with Global Feature Fusion.
4.2 Global Feature Fusion Module
@@
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers [54, 55, 56]. Among them, the Swin Transformer[54] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. By incorporating a sliding window approach similar to convolution, the Swin Transformer limits self-attention calculations within each sliding window, enabling local information interactions while reducing computational load and improving efficiency. The Swin Transformer also adopts a multi-scale hierarchical structure similar to that of CNNs, dividing the model into multiple stages. In each stage, a Patch Merging layer performs down-sampling to reduce the computational load, followed by the Swin Transformer Block, which applies self-attention within sliding windows. The Swin Transformer Block includes both Window Attention and Shifted Window Attention mechanisms; the former computes self-attention within each window, while the latter enables information exchange across windows.
@@
The Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy. Thus, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
4.3 Kernel Feature Fusion Module
Dynamic convolution is a specialized convolution operation that dynamically adjusts kernel parameters based on the input data. In conventional convolutions, kernel parameters remain fixed after training and are independent of the input data. In contrast, dynamic convolution adapts kernel parameters for each input, enhancing the network’s flexibility and generalization. For specific tasks, dynamic convolution often achieves better performance than standard convolution. CondInst[57] is an instance segmentation framework based on dynamic convolution, which generates convolution kernel parameters dynamically for each instance based on its features and applies them in the segmentation head to filter the instance’s mask. Since each kernel parameter only needs to handle a single instance, this approach reduces the network’s learning burden and convolutional computation load, while also eliminating the ROI operations required in methods like Mask R-CNN, creating a faster, more streamlined instance segmentation framework. Other frameworks, such as SOLO[58], SOLOv2[59], and CondLane[38] , have also applied dynamic convolution in instance segmentation, with CondLane specifically designed for lane instance segmentation, achieving robust performance with its dynamic convolution-based lane detection strategy.
Building on these studies, our lane detection head also incorporates dynamic convolution. This detection head includes two components: lane instance detection and lane mask segmentation. During instance detection, the network identifies lane instances within the image and generates a set of dynamic convolution kernel parameters for each lane. These kernel parameters are applied in the lane mask segmentation head to isolate each lane’s instance segmentation. In typical dynamic convolution-based instance segmentation frameworks, the center position features of each instance are used to generate the dynamic convolution kernels. However, for lane detection, this is less effective, as lanes are elongated structures, making it challenging for center features to capture information along the full lane length. Therefore, following CondLane’s approach, we use the starting position features of each lane to guide the dynamic convolution kernel generation. Specifically, we design an instance detection head in the third stage of the neck network’s feature map, comprising two branches: one branch predicts the starting points of lanes, generating a heatmap for these points; the other branch predicts a set of dynamic convolution kernels for each position, using the kernel at the lane’s starting position for subsequent segmentation.
However, the existing method using lane starting points does not account for the inherent structural and positional relationships among lane instances. In a top-down view, lanes are often parallel, similar in shape, and have consistent spacing. To fully leverage this structural information, we designed a Kernel Feature Fusion Module (KFFM) to enable the network to automatically learn and predict relationships between lane instances. The KFFM is shown in the bottom right of the Fig. 8.
4.4 Low-level Edge Enhanced Module
@@
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 10. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a $3\times 3$ and a $1\times 1$ convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 8. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
4.5 Loss functions
@@
4.5.1 Lane start point prediction
@@
$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
4.5.2 Row-based segmentation prediction
@@
4.5.3 Vertical range prediction
@@ 5.1 Datasets
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution $1640\times 590$ pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also $1640\times 590$ pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of $1280\times 720$ pixels.
5.2 Evaluation Metric
Lane detection tasks are typically regarded as a semantic segmentation problem; therefore, the evaluation metrics for lane detection are often similar to those used in segmentation tasks, specifically the confusion matrix: TP (True Positive), FP (False Positive), FN (False Negative), and TN (True Negative). In the context of lane detection, TP typically represents the correctly predicted lane line pixels, FP denotes incorrectly predicted lane points, FN refers to lane line points that were not detected by the model but are actually present, and TN indicates correctly predicted background pixels. Furthermore, the nature of lane detection tasks introduces some unique characteristics that differentiate their evaluation criteria from those of standard segmentation tasks.
For FoggyLane and FoggyCULane, we adopt the evaluation metrics of SCNN[14], connecting the labeled or predicted lane points sequentially to form a long line segment with a width of 30 pixels centered around the connecting line. We then calculate the Intersection over Union (IoU) between the predicted area and the ground truth area. If the IoU of the predicted area and the ground truth area exceeds a certain threshold (typically set at 0.5), the lane detection is considered correct; otherwise, it is deemed incorrect. Based on this, we calculate the Precision, Recall, and F1-score, with the following formulas:
@@ 5.3 Implementation Details
We conducted our experiments on the NVIDIA GeForce RTX 3090 platform. To achieve better and more stable performance[60], we first pre-trained the SwinGFFM backbone network on the COCO dataset for 500 epochs. Subsequently, during the training process for lane detection tasks, we loaded the pre-trained weights of the backbone network and resized the input images to 800 × 320. Additionally, we applied various data augmentation techniques to the input images, including rotation, horizontal flipping, brightness adjustment, and random scaling and cropping. During training, the batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay strategy. For detailed network parameters and training settings, please refer to our open-source code.
Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
For objective benchmarking against established lane detection methods including SCNN, RESA, LaneATT, CLRNet, FLAMNet, LSTR, DBNet, UFLD, and CondLane, we replicated these methods on our experimental platform. Given the differences in the architecture of each algorithm, the required batch size, number of training epochs, and learning rate vary, and forcing uniform training settings could negatively impact model performance. Therefore, we adhered to the training configurations provided in the respective papers. Specifically, as the FoggyCULane and FoggyTusimple datasets are extended versions of the CULane and Tusimple datasets, we adopted the original training parameters from CULane and Tusimple when evaluating the algorithms on FoggyCULane and FoggyTusimple. For the FoggyLane dataset, we used the training parameters from CULane and extended the training epochs to ensure sufficient training, thereby obtaining reliable test results.
5.4 Quantitative Results
5.4.1 Results on FoggyLane
@@
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
@@
From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
5.4.2 Results on FoggyCULane
@@
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
5.4.3 Results on FoggyTusimple
@@
As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
@@
Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
@@
5.5 Ablation Study
To explore the effectiveness of each part of the model, we conducted ablation studies on both real and synthetic datasets, namely FoggyLane and FoggyCULane. We have studied the roles of the GFFM, KFFM, and LEEM modules in the model, with Swin-t as the backbone network. As shown in the Tables. 5 and 6, adding each module boosts the model’s performance on both datasets. Among them, GFFM contributes the most to the model’s improvement, with increases of 1.07 on FoggyLane and 0.41 on FoggyCULane. This indicates that GFFM can effectively capture both local and global information, enriching the feature representation. Building on the foundation of GFFM, the addition of the KFFM module further increases the F1-score by 0.63 on FoggyLane and by 0.26 on FoggyCULane. GFFM is capable of automatically learning the associations between lane instance instances, making full use of relevant information. On this basis, the introduction of the LEEM module enhances the network’s focus on edge information. The F1-score increases by 0.28 on FoggyLane and by 0.14 on FoggyCULane.
Table 5: Ablation experiments of different components on the FoggyLane dataset.
@@ 5.5 Ablation Study
Table 6: Ablation experiments of different components on the FoggyCULane dataset.
@@ 5.6 Model deployment
After considering the deployment of several edge computing tasks[61, 62], we further evaluate the performance of our algorithm in real-world scenarios. In addition to experiments on existing datasets, we also deployed and tested the model on the vehicle. Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig. 14. This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit. The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
6 Conclusion
@@ Acknowledgments
Compliance with ethics guidelines
Ronghui Zhang, Yuhang Ma, Tengfei Li, Ziyu Lin, Yueying Wu, Junzhou Chen, Lin Zhang, Jia Hu,Tony Z. Qiu and Konghui Guo declare that they have no conflict of interest or financial conflicts to disclose.

### v5

@@ IV-A Overall structure
Figure 6: Overview of the proposed network.
Figure 7: Backbone Network with Global Feature Fusion. The Swin Transformer’s window attention mechanism mainly focuses on local features. Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes. To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
IV-B Global Feature Fusion Module
@@
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers[52, 53, 54]. Among them, the Swin Transformer[52] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 7, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
IV-C Kernel Feature Fusion Module
Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution. CondInst[55] is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency. Similarly, CondLane[39] applies dynamic convolution for lane instance segmentation, achieving strong performance.
Building on this, our lane detection head also uses dynamic convolution, consisting of lane instance detection and mask segmentation. Unlike traditional methods, we generate dynamic convolution kernels using the lane’s starting position features rather than center features, which are more suitable for the elongated nature of lanes. We design an instance detection head with two branches: one predicts lane starting points and generates a heatmap, and the other generates dynamic convolution kernels for lane segmentation.
However, existing methods don’t fully utilize the structural relationships between lane instances. To address this, we introduce a Kernel Feature Fusion Module (KFFM) that enables the network to learn and predict relationships between lane instances, enhancing detection performance. KFFM is shown in the bottom right of Fig. 6.
IV-D Low-level Edge Enhanced Module
@@
Images in foggy conditions differ from those captured under good lighting primarily due to their blurriness and low contrast. This effect renders the edges of objects and details within the scene indistinct and hard to detect, impacting the algorithm’s ability to understand and recognize the scene accurately. Edge information, beyond simply marking object boundaries, also implicitly contains structural and spatial information about objects, forming a critical foundation for tasks such as object detection, segmentation, and tracking[tim1]. Using the Canny edge detection algorithm with identical threshold settings, we processed both foggy and normal images, as shown in the Fig. 8. It is evident that fog significantly degrades edge information, particularly affecting lane line edges, which become nearly indiscernible. Given that lane detection relies heavily on edge information—lane positioning is typically determined by the boundary between lane edges and the road background—both global scene information and detailed edge information are essential for accurate lane detection.
@@
TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
Specifically, we add an auxiliary path in the network to learn the edge features of the input image. First, we apply the Canny operator to the input image for edge detection, supplemented by lane line labels to reconstruct any missing lane edges, generating the training labels for the edge enhancement module. To avoid interfering with high-level semantic extraction and subsequent detection segmentation, while strengthening the low-level network’s learning of edge textures and other low-dimensional features, we introduce a branch at the second stage of the backbone network. This branch sequentially applies a 3 × 3 and a 1 × 1 convolution. We then compute the cross-entropy loss between the branch output and the edge labels. The detailed structure is illustrated in the bottom-left of the Fig. 6. Notably, this auxiliary enhancement path is only active during training, thus not increasing computation or processing time during deployment.
TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
IV-E Loss functions
@@
IV-E1 Lane start point prediction
@@
$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$
IV-E2 Row-based segmentation prediction
@@
IV-E3 Vertical range prediction
@@ V-A Datasets
FoggyLane is a lane detection dataset specifically designed for foggy scenarios, containing 1,423 frames of foggy lane images for training and testing. All images are of resolution 1640x590 pixels.
FoggyCULane is synthesized from the large-scale challenging lane detection dataset CULane using a foggy image modeling approach. It maintains the same number of images as CULane, which includes 88,880 images for training and 34,680 images for testing. The resolution of all images is also 1640x590 pixels.
FoggyTusimple is synthesized from one of the most widely used benchmark datasets for lane detection, Tusimple, through the foggy image modeling method. At the same time, the dataset follows the same structure as Tusimple, comprising 3,268 training images, 358 validation samples, and 2,782 test cases. All visual data maintains a consistent resolution of 1280×720 pixels.
V-B Evaluation Metric
Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN). However, due to its unique structure, lane detection also adopts specialized metrics. Following SCNN’s[8] protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas. A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5. Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
@@ V-C Implementation Details
We conducted experiments on the NVIDIA GeForce RTX 3090 platform. The SwinGFFM backbone network was pre-trained on the COCO dataset for 500 epochs. During lane detection training, we loaded the pre-trained weights and resized input images to 800 × 320. Various data augmentation techniques, including rotation, flipping, brightness adjustment, and random scaling, were applied. The batch size was set to 16, with an initial learning rate of $3\times 10^{-4}$, using the Adam optimizer and a Step learning rate decay. For full network parameters and training details, please refer to our open-source code.
To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
V-D Quantitative Results
V-D1 Results on FoggyLane
@@
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
@@
From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
V-D2 Results on FoggyCULane
@@
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.81,which achieves a 0.81% improvement over the best alternative method.
V-D3 Results on FoggyTusimple
@@
As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
@@
TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
@@
V-E Ablation Study
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone. As shown in Tables V and VI, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
TABLE V: Ablation experiments of different components on the FoggyLane dataset.
@@ V-E Ablation Study
TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.
@@ V-F Model deployment
To further evaluate the performance of our algorithm in real-world scenarios, in addition to experiments on existing datasets, we also deployed and tested the model on the vehicle. Specifically, we implemented an edge computing system on the vehicle, as illustrated in the Fig. 12. This system comprises the NVIDIA Jetson AGX Orin edge computing device, a monitor, a camera, and a power supply unit. The results of the evaluation demonstrate that our approach can operate in real-time with high accuracy on the vehicle’s edge computing platform, despite the constraints on computational resources.
VI Conclusion